Right now, Susan's age is 4 more than four times Jane's age. In 20 years, Jane will be half of Susan's age then. How old are they now?
Answer by JulietG(1812)

S = 4J + 4 [Susan's age is 4 more than four times Jane's age.]
.5(S + 20) = J + 20 [In 20 years, Jane will be half of Susan's age then]
Substitute the known value of S from the first equation into the second.
.5(4J + 4 + 20) = J + 20
Add and Distribute
2J + 12 = J + 20
Subtract J from each side
J + 12 = 20
subtract 12 from each side
J = 8
.
Let's put it into the equation to prove.
Susan is 4 more than 4*Jane, so she is 36.
In 20 years, Susan will be 56. Jane will be 28
Is 28 half of 56? Yes.
Success!